Particles A and B carrying charge 3.0nC are at the base corners of an equilateral triangle 2.8 m on a side

How much work is required to bring a positively charged particle carrying a charge of 5.0nC from infinity to the apex if particle B is replaced by a particle carrying a charge of - 3.0nC and A and B are fixed?
Then
Calculate the electric potential energy for the new triangular distribution?
1
Expert's answer
2021-10-14T11:31:31-0400

Calculate the potential at the apex from charges: of equal charge but opposite sign:


"V_1=-V_2=|kq/r|,\\
V=V_1+V_2=0."

Determine the work:


"W=VQ=0."

The energy of the new triangular combination at the center of the triangle is


"U=\frac{2\sqrt3·kq_1q_2}{r}+\frac{2\sqrt3·kq_1q_3}{r}+\frac{2\sqrt3·kq_3q_2}{r},\\\space\\
U=-1·10^{-7}\text{ J}."